1. Are the following sequences well ordered as € +0? If not, arrange them so that they are. (a) Mn = (sinh(6/2)]2n for n = 0,(a) yes (b) no (c) no (d) yesanswers provided
Answer #1

. . Given sequence I k = (sinh (%), ha), 2 : taking & o we get Lemot = lim (sinh (J] - ( T+A (8)” . = (sfhh Oh =0 This sequent = ε INCE) &= Eloge taking eso lim (+3) = lim Eleje (ox form by. L Hospital scule = lema Ees limen = lenget = L sin ed} dem ts - lim 16 E-70 bf254@yo--) [éno = o-gpt este Elom (2nd et---> .. Eto do= In (E) along the re comme The© ¢ = -1-€ dim ei a lam E* -* -€ : - e-nos € = 6 - 6 = -1-8-48 lima = lun (e-l) E70 . E- lim t = lim (e-le-1 & 76--0- 5,0 1-1до -3 е-то = e-. = о А и